### v1

# Wasserstein Evolution : Evolutionary Optimization as Phase Transition

Abstract
This paper establishes a novel connection between evolutionary computation and statistical physics by formalizing evolutionary optimization as a phase transition process. We introduce Wasserstein Evolution (WE), a principled optimization framework that implements the Wasserstein gradient flow of a free energy functional, mathematically bridging evolutionary dynamics with thermodynamics. WE directly translates the physical competition between potential gradient forces (exploitation) and entropic forces (exploration) into algorithmic dynamics, providing an adaptive, theoretically grounded mechanism for balancing exploration and exploitation. Experiments on challenging benchmark functions demonstrate that WE achieves competitive convergence performance while maintaining dramatically higher population diversity than classical methods (GA, DE, CMA-ES).This superior entropy preservation enables effective navigation of multi-modal landscapes without premature convergence, validating the physical interpretation of optimization as a disorder-to-order transition. Our work provides not only an effective optimization algorithm but also a new paradigm for understanding evolutionary computation through statistical physics.
Keywords: Wasserstein Evolution, Phase Transition, Free Energy Minimization, Evolutionary Optimization, Statistical Physics
1  Introduction
In real-world complex adaptive systems, simple local rules can spontaneously drive the system from macroscopic disorder to emergent order, achieving a collective functionality where the whole exceeds the sum of its parts. In nature, the process of biological evolution, facilitated by genetic crossover, mutation, and natural selection, shifts the genotypic distribution from regions of low fitness to those of higher fitness, a phenomenon well-modeled by evolutionary computation[1]. In physical systems, non-equilibrium thermodynamic processes, such as glass transitions, are driven from disorder toward order through free energy minimization, with statistical physics providing the theoretical formalization of these mechanisms[2]. In recent years, artificial intelligence models, particularly neural networks, demonstrate how simple connectivity rules among neurons, together with backpropagation, enable the learning of underlying data distributions for accurate prediction[3].
From an optimization perspective, evolutionary computation minimizes given objective functions by introducing the concept of fitness. Statistical physics models, in turn, drive free energy minimization spontaneously through the interplay of potential fields and entropy. Neural networks, meanwhile, fit given data by optimizing loss functions. Many studies have explored the relationship between learning and evolution, which computationally corresponds to the connection between evolutionary computation and neural networks[4]. On one hand, evolutionary algorithms serve as black‑box optimizers that naturally optimize hyperparameters and architectures of neural networks—a field known as neuroevolution[5]. Research has also revealed profound links between neuroevolution and gradient‑based optimization[6]. On the other hand, deep learning models, particularly neural networks, can act as surrogate models to accelerate the update process in evolutionary algorithms[7], guide the design and selection of evolutionary operators[8], or even inspire entirely new algorithms[9]. Beyond these connections, significant research has investigated the deep ties between statistical physics and neural networks. Early neural models—such as Hopfield networks[10] and Boltzmann machines—were directly inspired by statistical physics models[11]. Moreover, in recent years, theories grounded in statistical physics concepts have been developed to explain the generalization and convergence behaviors of deep learning, for instance, linking neural scaling laws to phase transition theory[12] and interpreting grokking phenomena through statistical‑physical frameworks[13].
Fig. 1: Loss Function Landscape, Fitness Landscape, and Energy Landscape in Learning, Evolution, and Physical Systems
However, although extensive research has elucidated the deep connections between learning and evolution as well as between learning and physics, comparatively little attention has been paid to the relationship between evolutionary computation and statistical physics.The convergence process in evolutionary algorithms inherently involves a dynamic shift between exploration and exploitation[14].In the initial phase, the algorithm broadly explores the solution space to identify promising regions; later, it intensively exploits these areas to refine the solution. This transition mirrors the competition observed in free energy minimization in physical systems between the potential gradient force and the entropic force[15].The potential gradient guides the system toward lower energy states, while the entropic force drives it toward disorder, corresponding to entropy increase.Formally, the exploitation phase in evolutionary algorithms appears analogous to the role of the potential gradient force, whereas exploration resembles the entropic force. Moreover, the transition from exploration to exploitation during convergence seems to parallel the shifting dominance between these two forces as the system cools. Furthermore, the evolution of the population from initial disorder toward ordered, high-fitness regions resembles a disorder-to-order transition, akin to phase transitions such as a liquid solidifying into a solid[16].This series of correspondences raises a fundamental question: are the parallels between evolutionary computation and statistical phase transitions merely superficial, or do they reflect a deeper mathematical duality?
Fig. 2: The convergence of evolutionary optimization is like liquid solidification.
Inspired by this question, this study proposes Wasserstein Evolution (WE), a novel optimization framework that bridges evolutionary computation and statistical physics through the principles of Wasserstein gradient flow and free energy minimization. Experimental results demonstrate that, compared to classical evolutionary methods such as Genetic Algorithm (GA)[17], Differential Evolution (DE)[18], and Covariance Matrix Adaptation Evolution Strategy (CMA-ES)[19], WE achieves excellent convergence performance on a suite of benchmark functions including Rastrigin, Beale, Himmelblau, Six-Hump Camel Back, and Holder Table. Notably, WE maintains highest entropy upon convergence, enabling it to better capture multimodal solutions and avoid premature convergence. The dynamic evolution of population entropy further reflects a phase-transition-like progression from disorder to order, consistent with the theoretical underpinnings of the algorithm.The main contributions of this paper are summarized as follows:
  1. 1.
    We establish a formal correspondence between evolutionary optimization and statistical phase transition theory, proposing a free-energy-minimization-based optimization paradigm grounded in Wasserstein gradient flows.
  2. 2.
    We design the Wasserstein Evolution, which adaptively balances exploration and exploitation by translating the competition between potential gradient forces and entropic forces—central to physical systems—into an evolutionary optimization context.
  3. 3.
    We conduct extensive experiments on classical multimodal benchmark functions, systematically validating WE’s performance advantages—particularly its ability to sustain population diversity—and offering a new solution approach for complex optimization problems.
  4. 4.
    We provide novel interdisciplinary insights that reveal deep connections between optimization algorithms and physical systems, advancing the theoretical understanding of evolutionary convergence mechanisms and opening new pathways for applying statistical physics concepts in computational intelligence.
The central insight of this work is that the convergence of evolutionary algorithms can be understood as a statistical physical phase transition, with Wasserstein gradient flow offering a rigorous mathematical description of this process. This perspective not only deepens our understanding of optimization algorithms but also provides a new theoretical foundation for designing more efficient and robust intelligent optimization methods.
The remainder of this paper is organized as follows. Section 2 introduces the fundamental background of evolutionary algorithms and phase transitions. Section 3 elaborates on how evolutionary optimization corresponds to the phase transition process and presents the mathematical model of WE. Section 4 presents and analyzes the performance comparison between WE and competing algorithms on benchmark test functions. Finally, Section 5 concludes the paper and outlines potential directions for future research.
2  Background
2.1  Evolutionary Algorithms
Evolutionary algorithms (EAs) have developed as a family of population-based metaheuristic optimization methods inspired by biological evolution. The earliest and most classical form of EA can be traced back to the Genetic Algorithm (GA), introduced by Holland in 1992[17]. GA mimics natural selection by utilizing operators such as selection, crossover, and mutation to evolve a population of candidate solutions toward optimality. Subsequently, the EA framework has expanded to include several major variants, notably Differential Evolution (DE)[18] and Evolution Strategy (ES)[20]. Differential Evolution operates through vector-based mutation and crossover operations, typically employing strategies like DE/rand/1 or DE/best/1 to generate new trial vectors. Representative algorithms include DE/rand/1/bin and its adaptive successors such as JADE[21] and SHADE[22]. In contrast, Evolution Strategy emphasizes self-adaptive mutation and deterministic selection, often modeling mutation step sizes via covariance matrices, as exemplified by the well-known Covariance Matrix Adaptation Evolution Strategy (CMA-ES)[19].
Despite these developments—from classical Genetic Algorithms to Differential Evolution, Evolution Strategy, and modern variants—the core of EA design consistently revolves around balancing and transitioning between exploration and exploitation[14]. In the early iterations, algorithms prioritize exploration to broadly sample the search space and avoid premature convergence. As the evolution progresses, the focus gradually shifts toward exploitation to refine promising regions and converge to high-quality solutions. This adaptive trade-off between global search and local refinement remains a foundational principle in evolutionary computation.
Fig. 3: The evolution of humanity
2.2  Phase Transition
Phase transitions in physics describe fundamental transformations between different states of matter, with common examples including liquid–solid transitions (freezing)[23], gas–liquid transitions (condensation)[24], and magnetic transitions[25].A particularly illustrative example is the solidification process of a liquid upon cooling. During this phase transition, the system is driven by the interplay of two competing forces: the potential energy gradient, which favors the ordered, low-energy solid state, and the entropic force, which promotes the disordered, high-entropy liquid state. The potential gradient drives the system toward configurations with lower energy, while entropy resists complete ordering by maintaining thermal fluctuations and configurational diversity[26, 27].
As temperature decreases, the balance shifts in favor of the potential term, and the system undergoes a transition from a disordered liquid phase to an ordered solid phase. This progression from disorder to order bears a striking resemblance to the exploration–exploitation transition in evolutionary algorithms. In the early stages of evolution, high population entropy corresponds to broad exploration of the search space (analogous to the disordered liquid phase). Over time, selection pressure increases, entropy decreases, and the population converges toward optimal solutions (analogous to the ordered solid phase). Thus, the thermodynamics of phase transitions provides a natural physical metaphor for understanding how evolutionary algorithms dynamically shift from exploration to exploitation during optimization.
Fig. 4: Glass Transition in Liquid-Solid Phase Transition[28]
3  Evolutionary Optimization as Phase Transition
3.1  Wasserstein Gradient Flow
The Wasserstein gradient flow provides a rigorous geometric framework for describing the dynamical evolution of probability distributions. It generalizes the classical notion of gradient descent from Euclidean space to the space of probability measures endowed with the Wasserstein metric[29, 30, 31].
Given a functional $\mathcal{F}[\rho]$ defined on the space of probability densities $\rho(\mathbf{x})$, the Wasserstein gradient flow describes how $\rho$ evolves in time to minimize $\mathcal{F}$. The evolution is governed by the continuity equation:
$\frac{\partial\rho_{t}}{\partial t}=\nabla\cdot\left(\rho_{t}\nabla\frac{\delta\mathcal{F}[\rho_{t}]}{\delta\rho}\right),\qquad t\geq 0,$ (1)
where $\frac{\delta\mathcal{F}}{\delta\rho}$ denotes the first variation of the functional, and $\rho_{t}(\mathbf{x})=\rho(\mathbf{x},t)$ represents the probability density at time $t$.
In the context of evolutionary optimization, we consider the following free energy functional, which naturally incorporates both exploitation (fitness optimization) and exploration (entropy maximization):
$\mathcal{F}_{\beta}[\rho]=\underbrace{\mathbb{E}_{\mathbf{x}\sim\rho}[f(\mathbf{x})]}_{\text{Potential Energy (Exploitation)}}-\frac{1}{\beta}\underbrace{S[\rho]}_{\text{Entropy (Exploration)}},$ (2)
where $f(\mathbf{x})$ is the fitness function, $S[\rho]=-\int\rho(\mathbf{x})\log\rho(\mathbf{x})\,d\mathbf{x}$ is the Shannon entropy, and $\beta>0$ is the inverse temperature parameter controlling the trade-off between the two competing terms.
The Wasserstein gradient flow for this free energy functional leads to the following particle-level dynamics. Let $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}$ be $N$ particles approximating the distribution $\rho_{t}$. Their trajectories follow the mean-field equation:
$\frac{d\mathbf{x}_{i}}{dt}=-\nabla f(\mathbf{x}_{i})-\frac{1}{\beta}\nabla\log\rho_{t}(\mathbf{x}_{i}),\qquad i=1,\dots,N.$ (3)
The first term $\nabla f(\mathbf{x}_{i})$ represents the potential gradient force, which drives particles toward regions of higher fitness (exploitation). The second term $\frac{1}{\beta}\nabla\log\rho_{t}(\mathbf{x}_{i})$ represents the entropic force (or diversity pressure), which pushes particles away from high-density regions toward less explored areas (exploration).
The connection to physical phase transitions becomes evident through the temperature parameter $\beta$. At high temperature ($\beta\to 0$), the entropic term $\frac{1}{\beta}\nabla\log\rho$ dominates, resulting in a disordered phase characterized by broad exploration. As the system cools ($\beta$ increases), the potential gradient term becomes increasingly dominant, leading to an ordered phase focused on local exploitation. This continuous shift from exploration to exploitation mirrors the liquid-to-solid phase transition in physical systems.
The Wasserstein metric ensures that the flow follows the path of steepest descent with respect to the intrinsic geometry of probability distributions, making it a natural framework for describing population dynamics in optimization.
3.2  Fokker-Planck Equation
The mathematical formalism of Wasserstein gradient flow establishes a profound connection with the Fokker-Planck equation from statistical physics, revealing that these two seemingly distinct formulations describe the same underlying physical process from complementary perspectives. This connection provides crucial insights into the thermodynamic interpretation of evolutionary optimization[32].
For the specific free energy functional $\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f]-\frac{1}{\beta}S[\rho]$, we compute its first functional derivative:
$\frac{\delta\mathcal{F}_{\beta}[\rho]}{\delta\rho}=f(\mathbf{x})+\frac{1}{\beta}(\log\rho(\mathbf{x})+1).$ (4)
Substituting this expression into the Wasserstein gradient flow equation yields:
$\frac{\partial\rho_{t}}{\partial t}$ $=\nabla\cdot\left[\rho_{t}\nabla\left(f(\mathbf{x})+\frac{1}{\beta}(\log\rho_{t}(\mathbf{x})+1)\right)\right]$
$=\nabla\cdot\left[\rho_{t}\nabla f(\mathbf{x})\right]+\frac{1}{\beta}\nabla\cdot\left[\rho_{t}\nabla\log\rho_{t}(\mathbf{x})\right]$
$=\nabla\cdot\left[\rho_{t}\nabla f(\mathbf{x})\right]+\frac{1}{\beta}\nabla^{2}\rho_{t}(\mathbf{x}),$ (5)
where the final equality follows from the identity $\rho_{t}\nabla\log\rho_{t}=\nabla\rho_{t}$. Equation (5) is precisely the Fokker-Planck (or Smoluchowski) equation:
$\frac{\partial\rho_{t}}{\partial t}=\nabla\cdot\left[\rho_{t}\nabla f(\mathbf{x})\right]+D\nabla^{2}\rho_{t},$ (6)
with diffusion coefficient $D=\frac{1}{\beta}$.
This Fokker-Planck equation describes the probability density evolution for particles undergoing overdamped Langevin dynamics:
$d\mathbf{x}_{t}=-\nabla f(\mathbf{x}_{t})dt+\sqrt{2D}d\mathbf{W}_{t},$ (7)
where $\mathbf{W}_{t}$ is a standard Wiener process. The mathematical equivalence can be formally established through Itô’s formula: for any smooth test function $\phi(\mathbf{x})$, applying Itô’s lemma to $\phi(\mathbf{x}_{t})$ and taking expectations yields the Kolmogorov backward equation, whose dual is precisely the Fokker-Planck equation (6).
The inverse temperature parameter $\beta$ plays a dual role in this correspondence. In the Wasserstein formulation, $\beta^{-1}$ weights the relative strength of the entropic force $\frac{1}{\beta}\nabla\log\rho_{t}$ against the potential gradient force $\nabla f$. In the Fokker-Planck formulation, $\beta^{-1}=D$ represents the diffusion coefficient controlling the intensity of stochastic fluctuations.
The stationary solution of the Fokker-Planck equation (6) is the Gibbs distribution:
$\rho_{\infty}(\mathbf{x})=\frac{1}{Z}e^{-\beta f(\mathbf{x})},\qquad Z=\int_{\Omega}e^{-\beta f(\mathbf{x})}d\mathbf{x},$ (8)
which minimizes the free energy $\mathcal{F}_{\beta}[\rho]$. This establishes a direct link between the long-time behavior of the stochastic dynamics and the equilibrium statistical mechanics of the system.
The convergence properties can be analyzed through the free energy dissipation. Along solutions of the Fokker-Planck equation, the free energy decreases monotonically:
$\frac{d}{dt}\mathcal{F}_{\beta}[\rho_{t}]$ $=\int_{\Omega}\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\frac{\partial\rho_{t}}{\partial t}d\mathbf{x}$
$=\int_{\Omega}\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\nabla\cdot\left[\rho_{t}\nabla\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\right]d\mathbf{x}$
$=-\int_{\Omega}\rho_{t}\left\|\nabla\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\right\|^{2}d\mathbf{x}\leq 0,$ (9)
where the final equality follows from integration by parts. This dissipation inequality quantifies the rate at which the system approaches equilibrium and provides a natural convergence criterion for optimization algorithms.
This mathematical equivalence reveals that evolutionary optimization can be viewed through two complementary lenses: as a deterministic gradient flow in the Wasserstein geometry of probability distributions, or as a stochastic process described by the Fokker-Planck equation. The former emphasizes the collective behavior of the population as a whole, while the latter focuses on the individual particle dynamics. Both perspectives converge to the same physical reality: the minimization of free energy through the interplay of deterministic forces (exploitation) and stochastic fluctuations (exploration), with the inverse temperature $\beta$ governing the transition from disordered exploration to ordered exploitation.
3.3  Free Energy Minimization
The principle of free energy minimization provides a unified perspective on the convergence behavior of both physical systems and optimization algorithms[33]. In statistical physics, the equilibrium state of a system at temperature $T=1/\beta$ is given by the Gibbs distribution that minimizes the Helmholtz free energy:
$\rho_{\text{eq}}(\mathbf{x})=\frac{1}{Z}e^{-\beta f(\mathbf{x})},\qquad Z=\int e^{-\beta f(\mathbf{x})}d\mathbf{x},$ (10)
where $Z$ is the partition function. This distribution minimizes the functional:
$\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f]-\frac{1}{\beta}S[\rho].$ (11)
The minimization process exhibits distinctive characteristics at different temperature regimes:
  • High temperature limit ($\beta\to 0$): The entropy term dominates, $\rho_{\text{eq}}(\mathbf{x})$ approaches a uniform distribution, maximizing exploration.
  • Low temperature limit ($\beta\to\infty$): The energy term dominates, $\rho_{\text{eq}}(\mathbf{x})$ concentrates around global minima of $f(\mathbf{x})$, maximizing exploitation.
  • Finite temperature: The distribution balances between exploration (high entropy) and exploitation (low energy).
The dynamical approach to equilibrium follows the Wasserstein gradient flow described in the previous section. This flow can be interpreted as a gradient descent in the space of probability distributions:
$\frac{d}{dt}\mathcal{F}_{\beta}[\rho_{t}]=-\int\rho_{t}(\mathbf{x})\left\|\nabla\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\right\|^{2}d\mathbf{x}\leq 0,$ (12)
which guarantees monotonic decrease of free energy over time.
In evolutionary computation, the population distribution $\rho_{t}(\mathbf{x})$ evolves to minimize a similar objective. The fitness function $f(\mathbf{x})$ corresponds to the energy landscape, while population diversity corresponds to entropy. The selection pressure parameter in evolutionary algorithms plays a role analogous to the inverse temperature $\beta$. As the algorithm progresses, increasing selection pressure corresponds to decreasing temperature, driving the population from a disordered, exploratory phase to an ordered, exploitative phase.
This formal correspondence suggests that evolutionary optimization can be viewed as a non-equilibrium thermodynamic process, where the population undergoes a phase transition from high-entropy exploration to low-entropy exploitation. The competition between selection pressure (driving convergence) and mutation/crossover operators (maintaining diversity) parallels the competition between energy minimization and entropy maximization in physical systems.
The mathematical framework of free energy minimization provides a principled way to design optimization algorithms that naturally balance exploration and exploitation through temperature scheduling, much like simulated annealing but operating at the population distribution level rather than individual solution level.
3.4  Wasserstein Evolution
Building upon the theoretical foundations of Wasserstein gradient flow and free energy minimization, we propose the Wasserstein Evolution (WE), which explicitly formulates evolutionary optimization as a phase transition process. WE directly implements the mean-field equation derived from the Wasserstein gradient flow of the free energy functional:
Algorithm 1 Wasserstein Evolution (WE)
1:Initialize population $\{\mathbf{x}_{i}^{(0)}\}_{i=1}^{N}\sim\text{Uniform}(\Omega)$
2:for $k=0,1,\dots,K-1$ do
3:  Compute current inverse temperature: $\beta_{k}=\beta_{0}+(\beta_{\text{final}}-\beta_{0})\cdot(k/K)^{2}$
4:  Estimate density: $\hat{\rho}_{k}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}K_{h}(\mathbf{x}-\mathbf{x}_{i}^{(k)})$
5:  for $i=1$ to NN do
6:   Compute fitness gradient: $\nabla f(\mathbf{x}_{i}^{(k)})\approx\frac{f(\mathbf{x}_{i}^{(k)}+\epsilon)-f(\mathbf{x}_{i}^{(k)}-\epsilon)}{2\epsilon}$
7:   Compute entropic force: $\nabla\log\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})\approx\frac{\nabla\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})}{\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})+\delta}$
8:   Update position: $\mathbf{x}_{i}^{(k+1)}=\mathbf{x}_{i}^{(k)}+\eta_{k}\left[-\nabla f(\mathbf{x}_{i}^{(k)})-\frac{1}{\beta_{k}}\nabla\log\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})\right]$
9:   Apply boundary constraints: $\mathbf{x}_{i}^{(k+1)}\leftarrow\text{clip}(\mathbf{x}_{i}^{(k+1)},\mathbf{l},\mathbf{u})$
10:  end for
11:  Compute population statistics: entropy $S_{k}$, diversity $D_{k}$, best fitness $f_{k}^{*}$
12:end for
13:return Best solution $\mathbf{x}^{*}$ and corresponding fitness $f(\mathbf{x}^{*})$
The algorithm consists of three key components:
1. Density Estimation: At each iteration, we estimate the population density $\hat{\rho}_{k}(\mathbf{x})$ using kernel density estimation (KDE):
$\hat{\rho}_{k}(\mathbf{x})=\frac{1}{Nh^{d}}\sum_{i=1}^{N}K\left(\frac{\mathbf{x}-\mathbf{x}_{i}^{(k)}}{h}\right),$ (13)
where $K$ is a Gaussian kernel and $h$ is the bandwidth parameter. The gradient $\nabla\log\hat{\rho}_{k}(\mathbf{x})$ is then computed numerically using finite differences.
2. Temperature Scheduling: The inverse temperature $\beta_{k}$ follows an annealing schedule:
$\beta_{k}=\beta_{0}+(\beta_{\text{final}}-\beta_{0})\cdot\left(\frac{k}{K}\right)^{2},$ (14)
where $\beta_{0}$ is the initial inverse temperature (high exploration) and $\beta_{\text{final}}$ is the final inverse temperature (high exploitation). The quadratic scheduling ensures a smooth transition from exploration to exploitation, mimicking the cooling process in physical phase transitions.
3. Force Composition: Each particle experiences two competing forces:
$\mathbf{F}_{\text{total}}(\mathbf{x})=-\nabla f(\mathbf{x})-\frac{1}{\beta_{k}}\nabla\log\hat{\rho}_{k}(\mathbf{x}).$ (15)
The relative strength of these forces evolves over time, with the entropic force ($\frac{1}{\beta_{k}}\nabla\log\hat{\rho}_{k}$) dominating in early iterations (high temperature phase) and the potential gradient force ($\nabla f$) dominating in later iterations (low temperature phase).
Theoretical Properties:
Theorem 1 (Free Energy Decrease).
Under mild regularity conditions on $f$ and with appropriate learning rate $\eta_{k}$, the population distribution $\rho_{k}$ in WE satisfies:
$\mathcal{F}_{\beta_{k}}[\rho_{k+1}]\leq\mathcal{F}_{\beta_{k}}[\rho_{k}]-\frac{\eta_{k}}{2}\mathbb{E}_{\rho_{k}}\left[\|\mathbf{F}_{\text{total}}(\mathbf{x})\|^{2}\right]+O(\eta_{k}^{2}).$ (16)
Theorem 2 (Phase Transition Characterization).
Define the order parameter $m_{k}=\mathbb{E}_{\rho_{k}}[\|\mathbf{x}-\boldsymbol{\mu}_{k}\|^{2}]$ where $\boldsymbol{\mu}_{k}=\mathbb{E}_{\rho_{k}}[\mathbf{x}]$. As $\beta_{k}$ increases, WE exhibits:
  • Disordered phase ($\beta_{k}<\beta_{c}$): $m_{k}$ remains large, population entropy high
  • Critical point ($\beta_{k}\approx\beta_{c}$): Rapid decrease in $m_{k}$, entropy drops sharply
  • Ordered phase ($\beta_{k}>\beta_{c}$): $m_{k}$ small, population concentrated near optima
WE provides a principled approach to evolutionary optimization that explicitly models the exploration-exploitation trade-off as a physical phase transition. By directly implementing the Wasserstein gradient flow of free energy, WE maintains theoretical coherence while achieving practical optimization performance, as demonstrated in our experimental results.
4  Experiments
To empirically validate the theoretical framework and algorithmic performance of Wasserstein Evolution (WE), we conduct comprehensive experiments on five classical benchmark functions: Rastrigin, Beale, Himmelblau, Six-Hump Camel Back, and Holder Table. These functions represent diverse optimization challenges, including high-dimensional multimodality, symmetry, and deceptive landscapes. We compare WE against three well-established evolutionary algorithms: Genetic Algorithm (GA), Differential Evolution (DE), and Covariance Matrix Adaptation Evolution Strategy (CMA-ES).
4.1  Experimental Setup
We conducted experiments on five classical benchmark functions to evaluate the Wasserstein Evolution (WE) against three established evolutionary algorithms: Genetic Algorithm (GA), Differential Evolution (DE), and Covariance Matrix Adaptation Evolution Strategy (CMA-ES). These functions, detailed in Table 1, were selected for their diverse characteristics, including multimodality, symmetry, and deceptive landscapes, which provide a comprehensive testbed for assessing exploration-exploitation balance.
Table 1: Benchmark Functions Used in the Experiments
Name Formula Features
Rastrigin $f(\mathbf{x})=10n+\sum_{i=1}^{n}\left[x_{i}^{2}-10\cos(2\pi x_{i})\right]$, $\mathbf{x}\in[-5.12,5.12]^{n}$ ($n=2$). Highly multimodal, separable. Global optimum: $f(\mathbf{0})=0$.
Beale $f(x,y)=(1.5-x+xy)^{2}+(2.25-x+xy^{2})^{2}+(2.625-x+xy^{3})^{2}$, $x,y\in[-4.5,4.5]$. Non-convex, unimodal with a narrow valley. Global optimum: $f(3,0.5)=0$.
Himmelblau $f(x,y)=(x^{2}+y-11)^{2}+(x+y^{2}-7)^{2}$, $x,y\in[-6,6]$. Symmetric with four equal-valued global minima. Minima: $(3.0,2.0)$, (2.805,3.131)(-2.805,3.131), (3.779,3.283)(-3.779,-3.283), (3.584,1.848)(3.584,-1.848), $f^{*}=0$.
Six-Hump Camel Back $f(x,y)=\left(4-2.1x^{2}+\frac{x^{4}}{3}\right)x^{2}+xy+(-4+4y^{2})y^{2}$, $x\in[-3,3],y\in[-2,2]$. Two global minima and four local minima. Global minima: $f(0.0898,-0.7126)=f(-0.0898,0.7126)\approx-1.0316$.
Holder Table $f(x,y)=-\left|\sin(x)\cos(y)\exp\left(\left|1-\frac{\sqrt{x^{2}+y^{2}}}{\pi}\right|\right)\right|$, $x,y\in[-10,10]$. Highly multimodal with numerous local minima. Four global minima: $f(\pm 8.05502,\pm 9.66459)\approx-19.2085$.
All algorithms were configured with a population size of $N=30$ and executed for $K=500$ iterations. For the proposed WE algorithm, we set the initial inverse temperature $\beta_{0}=0.1$, the final inverse temperature $\beta_{\text{final}}=1000$, the base learning rate $\eta_{0}=0.1$ with a linear decay schedule, and the kernel density estimation (KDE) bandwidth $h=0.5$.
Performance is assessed using two complementary metrics. The first is the Potential Energy, defined as the best function value $f(\mathbf{x}^{*})$ found by the algorithm, which quantifies its exploitation success. The second is the Differential Entropy, which quantifies the population diversity. For a final population $\{\mathbf{x}_{i}\}_{i=1}^{N}$, the entropy $S$ is estimated via multivariate Gaussian Kernel Density Estimation (KDE):
$S=-\frac{1}{N}\sum_{i=1}^{N}\log\hat{\rho}(\mathbf{x}_{i}),$ (17)
where $\hat{\rho}$ is the KDE estimate. A higher entropy value indicates a more diverse, multi-modal population distribution, reflecting stronger exploration capacity.
4.2  Results and Analysis
The performance of all four algorithms across the five benchmark functions is summarized in Table 2. Each cell reports the final potential energy $U$ (corresponding to the optimal function value) and the corresponding differential entropy $S$.
Table 2: WE and Its Competitors’ Performance on Potential Energy $U$ and Entropy $S$
Function WE GA DE CMA-ES
Rastrigin $U$: 0.00, $S$: 3.57 $U$: 0.00, $S$: -34.15 $U$: 0.00, $S$: -34.26 $U$: 0.00, $S$: -34.63
Beale $U$: 0.00, $S$: -3.19 $U$: 0.02, $S$: -21.44 $U$: 0.00, $S$: -33.92 $U$: 0.00, $S$: -33.43
Himmelblau $U$: 0.00, $S$: 3.98 $U$: 0.06, $S$: -34.08 $U$: 0.00, $S$: -34.05 $U$: 0.00, $S$: -2.16
Six-Hump Camel Back $U$: -1.03, $S$: 1.54 $U$: -1.03, $S$: -34.23 $U$: -1.03, $S$: -34.23 $U$: -1.03, $S$: -3.38
Holder Table $U$: -19.21, $S$: 6.19 $U$: -19.21, $S$: -34.86 $U$: -19.14, $S$: -34.02 $U$: -19.18, $S$: 5.98
Fig. 5: Final Population Distribution of Different Algorithms
The results reveal several key insights:
1. Superior Entropy Preservation: WE consistently achieves differential entropy values that are orders of magnitude higher than GA and DE, and significantly higher than CMA-ES on most functions. This empirically validates WE’s core mechanism of maintaining exploration pressure through the entropic force term $\frac{1}{\beta}\nabla\log\rho_{t}(\mathbf{x})$.
2. Competitive Convergence: Despite maintaining high entropy (diversity), WE matches or outperforms the other algorithms in finding the global minimum (or a very good one) across all tested functions. This demonstrates that the Wasserstein gradient flow of free energy $\mathcal{F}_{\beta}[\rho]$ successfully navigates the trade-off between exploration and exploitation.
3. Theoretical Alignment with Boltzmann Distribution: The observed behavior is a direct consequence of the theoretical steady-state of the Wasserstein gradient flow. Minimizing the free energy functional $\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f]-\frac{1}{\beta}S[\rho]$ leads to the Gibbs-Boltzmann distribution:
$\rho_{\beta}(\mathbf{x})=\frac{1}{Z(\beta)}\exp(-\beta f(\mathbf{x})),\qquad Z(\beta)=\int\exp(-\beta f(\mathbf{x}))d\mathbf{x}.$ (18)
WE’s dynamics drive the population distribution $\rho_{t}$ towards this $\rho_{\beta}$, explaining why it naturally maintains diversity on multi-modal functions. The annealing schedule for $\beta_{k}$ controls the phase transition: at low $\beta$ (high temperature), $\rho_{\beta}$ is diffuse (high entropy, exploration); at high $\beta$ (low temperature), $\rho_{\beta}$ concentrates on minima (low entropy, exploitation). 4. Implications for Multi-modal Optimization: The significantly higher entropy values for WE indicate that its final population is not collapsed to a single point or a small cluster but remains distributed across multiple basins of attraction. This is particularly valuable for real-world optimization problems where identifying multiple good solutions (global or local) is often desired, rather than just a single optimum.
In conclusion, the experiments confirm that Wasserstein Evolution effectively operationalizes the theoretical concept of free energy minimization via gradient flow. It achieves a fundamentally different balance between exploration and exploitation compared to classical evolutionary algorithms, resulting in superior diversity preservation without compromising solution quality, precisely aligning with the behavior predicted by its underlying statistical physics framework.
5  Conclusion
This paper has established a novel theoretical and algorithmic framework that bridges evolutionary optimization with statistical physics through the lens of phase transition theory. We have demonstrated that the convergence dynamics of evolutionary algorithms can be rigorously formalized as a free energy minimization process via Wasserstein gradient flows. Building upon this foundation, we proposed Wasserstein Evolution (WE), an algorithm that explicitly models the exploration-exploitation trade-off as a competition between potential gradient forces and entropic forces, mirroring the disorder-to-order transition observed in physical systems.
The core contribution of this work is threefold. First, we established a formal mathematical correspondence between evolutionary optimization and statistical phase transitions, providing a unified theoretical perspective on population-based search algorithms. Second, we designed WE, an algorithm that implements the Wasserstein gradient flow of free energy, naturally balancing exploration and exploitation through temperature-controlled entropy dynamics. Third, extensive experiments on classical benchmark functions validated WE’s superior performance, particularly its ability to maintain significantly higher population diversity while achieving competitive or better convergence compared to established algorithms like GA, DE, and CMA-ES.
Looking forward, several promising research directions emerge from this work. First, future studies should extend the comparison of WE with more advanced evolutionary algorithms, including state-of-the-art variants like L-SHADE and neuroevolution-based approaches. Second, the application scope of WE can be significantly expanded to various real-world optimization problems. Notably, WE’s ability to maintain population diversity makes it particularly suitable for neural network training, where it could help escape local minima and discover more robust solutions. Similarly, in neural architecture search (NAS), WE’s exploration capabilities could facilitate discovering diverse and efficient architectures beyond those found by gradient-based NAS methods. Additionally, engineering optimization problems with complex, multi-modal landscapes—such as structural design, parameter tuning in control systems, and portfolio optimization—represent natural domains where WE’s balanced search strategy could yield substantial benefits.
Beyond these applied directions, theoretical investigations into the critical behavior of WE, particularly the characterization of phase transition points and scaling laws in high-dimensional optimization landscapes, would deepen our understanding of the algorithm’s fundamental properties. Furthermore, exploring connections between WE and other gradient flow-based methods in machine learning, such as Stein Variational Gradient Descent (SVGD) and their extensions, could lead to new hybrid algorithms with enhanced capabilities.
In conclusion, this work has laid the groundwork for viewing evolutionary optimization through the powerful lens of statistical physics. By framing optimization as a phase transition process, we have not only developed a novel algorithm with desirable properties but also opened new avenues for interdisciplinary research that connects computational intelligence with fundamental physical principles.